Alanine (HA) is a diprotic amino acid with Kal-4.53x 10-3 and Ka2-1.36 × 10.10. Determine the pH of the following solutions A) 0.260 M alanine hydrochloride (H2A c) Number pH- B) 0.260 M alanine (HA) Number pH- ID C) 0.260 M sodium alaninate(Na A) Number
Answer #1

a)0.260 M alanine hydrochloride

H2A   ---------------> H+ + HA-

0.260                        0          0

0.260 - x                   x           x

Ka1 = x^2 / 0.260 - x

4.53 x 10^-3 = x^2 / 0.260 – x

1.18 x 10^-3 -4.53 x 10^-3 x = x^2

Solve as quadratic equation.

x = 0.032160 M

[H+] = 0.032160 M

pH= -log [H+] = -log 0.032160

pH = 1.50

B)

HA    ------------->    H+    + A-

0.260                      0             0

0.260- x                   x              x

Ka2 = x^2 / 0.260- x

1.36 x 10^-10 = x^2 / 0.260 - x

x = 5.95 x 10^-6

[H+] = 5.95 x 10^-6M

pH = -log[H+] = -log (5.95 x 10^-6)

pH = 5.23

c)

A- + H2O ------------> HA   + OH-

0.260                             0         0

0.260 - x                         x           x

Kb1 = x^2 / 0.260 - x

10^-14 / 1.36 x 10^-10 = x^2 / 0.260 - x

7.35 x 10^-5 = x^2 / 0.260 - x

x = 4.37 x 10^-3

[OH-] = 4.37 x 10^-3 M

pOH= -log [OH-] = -log (4.37 x 10^-3 )

        = 2.36

We know that,

pH + pOH = 14

then pOH = 14- pH

=1 4- 2.36

pH = 11.64